OpenStudy (anonymous):

how to do x over 8 minus one fourth equals one half

OpenStudy (anonymous):

\[\frac{ x }{ 8 }-\frac{ 1 }{ 4 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

\[solving for x... \frac{ x }{ 8} = \frac{ 1 }{ 2 } + \frac{ 1 }{ 4 }\rightarrow x=\frac{ 3 }{ 4 } * \frac{ 1 }{ 8 }\rightarrow x=\frac{ 3 }{ 32 }\]